OpenStudy (anonymous):

Can someone help me with this equation? Find f(a-2) when f(x)+5-x

OpenStudy (anonymous):

is it f(x) = 5 - x?

OpenStudy (anonymous):

yes sorry typo

OpenStudy (anonymous):

everywhere there is an x you plug in (a - 2) so f(a-2) = 5 - (a - 2) f(a - 2) = 5 - a + 2 f(a - 2) = 7 - a

OpenStudy (anonymous):

okay thank you so much i get it now :)